OpenStudy (friendzone):

Chris has $100 in his savings account and he adds $20 a month to the account. The total amount can be represented by the linear function T = 20x + 100. His goal is to have a total of $300 in four more months. What should he change in the function to reach this goal? A) Change the amount he adds each month to $50. B) Change the amount he adds each month to $40. C) Change the y-intercept of the linear function to $50. D) Change the y-intercepts of the linear function to $200.

OpenStudy (anonymous):

What do you think it is?

OpenStudy (anonymous):

Well, let's see in 4 months he gets: T = 20x + 100 T = 20(4) + 100 T = 80 + 100 T = 180 So we should probably change the amount he adds each month, so how much do you think it should be?

OpenStudy (anonymous):

40(3) = 120

OpenStudy (anonymous):

Let's try changing it to 50. T = 50x + 100 Plug in 4: T = 50(4) + 100 What's 50 * 4?

OpenStudy (friendzone):

50

OpenStudy (anonymous):

Yes, 50 is correct.

OpenStudy (anonymous):

50 * 4 = 200 200 + 100 = 300